7c+6+5c=234
Help!i need to solve for c

Answer by mathdoc314(58)


it is easy and fun to solve for c!
first you can combine all the c
12 c + 6 = 234
Move the 6 to the other side: subtract 6 from both sides
12 c = 228
Divide both sides by 12
c = 228/12 = 19
check if it is right:

7c + 6 + 5c =? 234
7*19 + 6 + 5*19 =? 234
133 + 6 + 95 = 234 YES!